# Note on shifted primes with large prime factors

### v1

@@ 2. Fundamental lemmas
where $P(z)=\prod_{\begin{subarray}{c}2<p\leq z\end{subarray}}p$ and $z\leq x^{1/2}$ is a parameter to be decided later. Then, we deduce from above notation that
$T^{\prime}_{c}(x)\leq\sum_{\begin{subarray}{c}a\in\mathcal{A}\\ a\in\mathcal{P}\end{subarray}}1\leq S(\mathcal{A},\mathcal{P},z)+\pi(z)=S(\mathcal{A},\mathcal{P},z)+O\big(x^{1/2}\big).$ (3.1)
For applications of Lemma 2.1, we now need to specify $\big|\mathcal{A}_{d}\big|$ and $r\big(\mathcal{A},d\big)$. Let $\varepsilon>0$ be an arbitrary small number and $D=x^{4/7-\varepsilon}$. For any $d\leq D$ with $d|P(z)$, we have
$\big|\mathcal{A}_{d}\big|=\sum_{\begin{subarray}{c}\ell p\leq x\\ \ell\leq x^{1-c},\,2|\ell\\ \ell p+1\equiv 0\!\!\!\!\!\pmod{d}\end{subarray}}1=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\ \sum_{\begin{subarray}{c}\ell p\leq x\\ \ell p\equiv-1\!\!\!\!\!\pmod{d}\end{subarray}}1=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\pi(x;\ell,-1,d).$
Next, we naturally approximate $\pi(x;\ell,-1,d)$ by ${\rm li}(x/\ell)/\varphi(d)$ and we write
$\big|\mathcal{A}_{d}\big|=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{{\rm li}(x/\ell)}{\varphi(d)}+r_{1}(\mathcal{A},d),$ (3.2)
where $r_{1}(\mathcal{A},d)$ is the error term:
$r_{1}(\mathcal{A},d)=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\bigg(\pi(x;\ell,-1,d)-\frac{{\rm li}(x/\ell)}{\varphi(d)}\bigg).$ (3.3)
Now we turn to estimate the sum over $\ell$ in (3.2), where the main term comes from.
$\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{{\rm li}(x/\ell)}{\varphi(d)}$ $=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{1}{\varphi(d)}\int_{2}^{x/\ell}\frac{{\rm{d}}t}{\log t}$
$=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{x/\ell}{\varphi(d)\log(x/\ell)}\Big\{1+O\Big(\frac{1}{\log x}\Big)\Big\}$
$=\Big\{1+O\Big(\frac{1}{\log x}\Big)\Big\}\frac{x}{\varphi(d)}\sum_{\begin{subarray}{c}\ell\leq x^{1-c}/2\\ (2\ell,\,d)=1\end{subarray}}\ \frac{1}{2\ell\log(x/\ell)}.$
The condition (2,d)=1(2,\,d)=1 is in fact redundant since $d|P(z)=\prod_{2<p\leq z}p$. To relax the condition $(\ell,\,d)=1$, we employ the Möbius inversion getting
$\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{{\rm li}(x/\ell)}{\varphi(d)}$ $=\Big\{1+O\Big(\frac{1}{\log x}\Big)\Big\}\frac{x}{2\varphi(d)}\sum_{\begin{subarray}{c}\ell\leq x^{1-c}/2\end{subarray}}\ \frac{1}{\ell\log(x/\ell)}\sum_{e\mid(\ell,\,d)}\mu(e)$
$=\Big\{1+O\Big(\frac{1}{\log x}\Big)\Big\}\frac{x}{2\varphi(d)}\sum_{e\mid d}\,\frac{\mu(e)}{e}\sum_{e\ell<x^{1-c}/2}\frac{1}{\ell\log(x/e\ell)}$
$=\Big\{1+O\Big(\frac{1}{\log x}\Big)\Big\}\frac{x}{2\varphi(d)}\big(S_{1}+S_{2}\big),$ (3.4)
@@
$S_{1}:=\sum_{\begin{subarray}{c}e\mid d\\ e<(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\sum_{\ell<x^{1-c}/2e}\frac{1}{\ell\log(x/e\ell)},\qquad S_{2}:=\sum_{\begin{subarray}{c}e\mid d\\ e\geq(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\sum_{\ell<x^{1-c}/2e}\frac{1}{\ell\log(x/e\ell)}.$
First, we estimate $S_{2}$ trivially
$S_{2}\ll\sum_{\begin{subarray}{c}e\mid d\\ e\geq(\log x)^{9}\end{subarray}}\frac{1}{e}\ll\frac{\tau(d)}{(\log x)^{9}},$ (3.5)
where $\tau(d)$ denotes the number of divisors of $d$.
And for $S_{1}$, we have by the partial summation
$S_{1}$ $=\sum_{\begin{subarray}{c}e\mid d\\ e<(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\bigg\{\int^{x^{1-c}/2e}_{1}\frac{{\rm{d}}t}{t\log(x/et)}+O\Big(\frac{1}{\log x}\Big)\bigg\}$
$=\sum_{\begin{subarray}{c}e\mid d\\ e<(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\bigg\{\log\bigg(\frac{\log(x/e)}{c\log x}\bigg)+O\Big(\frac{1}{\log x}\Big)\bigg\}$
$=\sum_{\begin{subarray}{c}e\mid d\\ e<(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\bigg\{\log\frac{1}{c}+O\bigg(\frac{\log\log x}{\log x}\bigg)\bigg\},$
where we have removed $\log e$ with an admissible error term in the last step thanks to the condition $e<(\log x)^{9}$. Now we reinsert the sum over $e$ with $e\geq(\log x)^{9}$ up to an error term as in (3.5) getting
$S_{1}$ $=\bigg(\sum_{e\mid d}\,\frac{\mu(e)}{e}-\sum_{\begin{subarray}{c}e\mid d\\ e\geq(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\bigg)\bigg\{\log\frac{1}{c}+O\bigg(\frac{\log\log x}{\log x}\bigg)\bigg\}$
$=\frac{\varphi(d)}{d}\bigg\{\log\frac{1}{c}+O\bigg(\frac{\log\log x}{\log x}\bigg)\bigg\}+O\bigg(\frac{\tau(d)}{(\log x)^{9}}\bigg).$ (3.6)
Combining (3), (3.5) and (3) we obtain
$\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{{\rm li}(x/\ell)}{\varphi(d)}=\frac{-\log c}{2}x\cdot\frac{1}{d}\Big\{1+O\Big(\frac{\log\log x}{\log x}\Big)\Big\}+O\bigg(\frac{x\tau(d)}{\varphi(d)(\log x)^{9}}\bigg),$
whence by (3.2) and (3.3) we arrive at
$|\mathcal{A}_{d}|=\frac{\omega(d)}{d}X+r(\mathcal{A},d)$
where
$X$ $=\frac{-\log c}{2}x\Big\{1+O\Big(\frac{\log\log x}{\log x}\Big)\Big\},\quad\omega(d)=1,$
and
$r(\mathcal{A},d)$ $=r_{1}(\mathcal{A},d)+O\bigg(\frac{x\tau(d)}{\varphi(d)(\log x)^{9}}\bigg)$
$=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\bigg(\pi(x;\ell,-1,d)-\frac{{\rm li}(x/\ell)}{\varphi(d)}\bigg)+O\bigg(\frac{x\tau(d)}{\varphi(d)(\log x)^{9}}\bigg).$
Now we are ready to apply Lemma 2.1. First we need to verify that the condition
$\prod_{\begin{subarray}{c}u\leq p<v\\ p\in\mathcal{P}\end{subarray}}\left(1-\frac{\omega(p)}{p}\right)^{-1}\leq\frac{\log v}{\log u}\left(1+\frac{K}{\log u}\right)$ (3.7)
holds for some absolute constant $K$. In fact, by Mertens’ formula it is easy to see
$\prod_{\begin{subarray}{c}u\leq p<v\\ p\in\mathcal{P}\end{subarray}}\left(1-\frac{\omega(p)}{p}\right)^{-1}=\prod_{\begin{subarray}{c}u\leq p<v\\ p>2\end{subarray}}\left(1-\frac{1}{p}\right)^{-1}=\frac{\log v}{\log u}\bigg(1+O\Big(\frac{1}{\log u}\Big)\bigg).$
Hence, with the choices of $X,\omega(d),r(\mathcal{A},d)$ as above, we deduce from Lemma 2.1 by taking $D=x^{4/7-\varepsilon}$ and $z=D^{\frac{1}{2}}=x^{\frac{2}{7}-\frac{\varepsilon}{2}}$ that
$S(\mathcal{A},\mathcal{P},z)$ $\leq XV(z)\bigg(F\Big(\frac{\log D}{\log z}\Big)+E\bigg)+\sum_{h<\exp(8/\varepsilon^{2})}\sum_{d|P(z)}\lambda^{+}_{h}(d)r\big(\mathcal{A},d\big),$
$=\frac{-\log c}{2}x\prod_{\begin{subarray}{c}2<p\leq z\end{subarray}}\Big(1-\frac{1}{p}\Big)\Big(F(2)+o(1)\Big)+\sum_{h<\exp(8/\varepsilon^{2})}\sum_{d|P(z)}\lambda^{+}_{h}(d)r_{1}\big(\mathcal{A},d\big)$
$\qquad+\sum_{h<\exp(8/\varepsilon^{2})}\,\sum_{d|P(z),\,d\leq D}O\bigg(\frac{x\tau(d)}{\varphi(d)(\log x)^{9}}\bigg)$ (3.9)
$=:S_{M}+S_{E1}+S_{E2},$
say. For the main term $S_{M}$, employing again Mertens’ formula
$\prod_{\begin{subarray}{c}2<p\leq z\end{subarray}}\Big(1-\frac{1}{p}\Big)=\prod_{\begin{subarray}{c}2<p\leq x^{\frac{2}{7}-\frac{\varepsilon}{2}}\end{subarray}}\Big(1-\frac{1}{p}\Big)=\frac{2e^{-\gamma}+o(1)}{\log(x^{2/7-\varepsilon/2})}=\Big(7e^{-\gamma}+o(1)\Big)\frac{1}{\log x},$
and noting that $F(2)=e^{\gamma}$, we obtain
$S_{M}$ $=\Big(-\frac{7}{2}\log c+o(1)\Big)\frac{x}{\log x}.$ (3.10)
We are now in a position to apply Lemma 2.2 to estimate the first error term $S_{E1}$, provided the function $\lambda^{+}_{h}(d)$ is well factorable of level $D=x^{4/7-\varepsilon}$. By taking $L_{1}=1$ and $L_{2}=x^{1-c}$ in Lemma 2.2, we have
$S_{E1}$ $=\sum_{h<\exp(8/\varepsilon^{2})}\sum_{\begin{subarray}{c}d\leq x^{4/7-\varepsilon}\end{subarray}}\lambda^{+}_{h}(d)\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,(\ell,d)=1\end{subarray}}r_{1}\big(\mathcal{A},d\big)$
$=\sum_{h<\exp(8/\varepsilon^{2})}\sum_{\begin{subarray}{c}d\leq x^{4/7-\varepsilon}\end{subarray}}\lambda^{+}_{h}(d)\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,(\ell,d)=1\end{subarray}}\bigg(\pi(x;\ell,-1,d)-\frac{{\rm li}\big(x/\ell\big)}{\varphi(d)}\bigg)$
$\ll_{\varepsilon}\frac{x}{(\log x)^{A}}$ (3.11)
for any $A>0$, which is admissible.
For the error term $S_{E2}$, it is easy to see
$S_{E2}\ll\frac{x}{(\log x)^{9}}\sum_{d\leq x^{4/7-\varepsilon}}\frac{\tau(d)}{\varphi(d)}\ll\frac{x}{(\log x)^{6}},$ (3.12)
which is also admissible.
Then, inserting (3.10), (3) and (3.12) into (LABEL:SM+SE1+SE2), we arrive at
$S(\mathcal{A},\mathcal{P},z)\leq\Big(-\frac{7}{2}\log c+o(1)\Big)\frac{x}{\log x}.$
Finally, $T^{\prime}_{c}(x)$ is estimated from (3.1) that

### v2

@@ 2. Fundamental lemmas
where $z\leq x^{1/2}$ is a parameter to be decided later. Then, we deduce from above notation that
$T^{\prime}_{c}(x)\leq\sum_{\begin{subarray}{c}a\in\mathcal{A}\\ a\,\textrm{is\, prime}\end{subarray}}1\leq S(\mathcal{A},\mathcal{P},z)+\pi(z)=S(\mathcal{A},\mathcal{P},z)+O\big(x^{1/2}\big).$ (3.1)
In general, to estimate $S(\mathcal{A},\mathcal{P},z)$ by sieve methods, we need to compute $\mathcal{A}_{d}=\{a\in\mathcal{A}:d\mid a\}$ as the form
where $X$ is a convenient close approximation to $|\mathcal{A}|$ which should be independent of $d$, $\omega(d)$ is a multiplicative function and $r\big(\mathcal{A},d\big)$ is the error term. Unfortunately, in our question it seems we can not extract such a $X$ independent of $d$ from $\mathcal{A}_{d}$. To avoid this obstacle, we shall start from the fundamental inequality
$\lambda_{d}^{-}*\mathbf{1}\leqslant\mu*\mathbf{1}\leqslant\lambda_{d}^{+}*\mathbf{1}$
involving sieve weights $\lambda^{\pm}_{d}$, based on an idea of Fouvry and Tenenbaum [16]. That is, we will apply Lemma 2.1 to give an upper bound for $S(\mathcal{A},\mathcal{P},z)$.
First, by the Möbius inversion formula, we have
$S(\mathcal{A},\mathcal{P},z)=$ $\mathop{\sum\,\sum}_{\begin{subarray}{c}\ell p\leq x\\ \ell\leq x^{1-c},\,2|\ell\end{subarray}}\textbf{1}*\mu\big((\ell p+1,\,P(z))\big)$
$\leq$ $\mathop{\sum\,\sum}_{\begin{subarray}{c}\ell p\leq x\\ \ell\leq x^{1-c},\,2|\ell\end{subarray}}\ \sum_{\begin{subarray}{c}d|\ell p+1\\ d|P(z)\end{subarray}}\lambda^{+}_{d}$
$=$ $\sum_{\begin{subarray}{c}d|P(z)\end{subarray}}\lambda^{+}_{d}\mathop{\sum\,\sum}_{\begin{subarray}{c}\ell p\leq x\\ \ell\leq x^{1-c},\,2|\ell\\ \ell p+1\equiv 0\!\!\!\!\!\pmod{d}\end{subarray}}1.$
Next we approximate the inner sum over $p$ by $\pi(x/\ell)/\varphi(d)$ getting
$S(\mathcal{A},\mathcal{P},z)$ $\leq\sum_{\begin{subarray}{c}d|P(z)\end{subarray}}\lambda^{+}_{d}\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{\pi(x/\ell)}{\varphi(d)}+\sum_{\begin{subarray}{c}d|P(z)\end{subarray}}\lambda^{+}_{d}\,r\big(\mathcal{A},d\big)$
$=:M(\mathcal{A},\mathcal{P},z)+\sum_{\begin{subarray}{c}d|P(z)\end{subarray}}\lambda^{+}_{d}\,r\big(\mathcal{A},d\big),$ (3.2)
say, where $r(\mathcal{A},d)$ is the error term defined by
$r(\mathcal{A},d)=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\bigg(\pi(x;\ell,-1,d)-\frac{\pi(x/\ell)}{\varphi(d)}\bigg).$
Thanks to the arguments of Iwaniec [23, Theorems 1 and 4], the sum $\sum_{d}\lambda^{+}_{d}\,r(\mathcal{A},d)$ can be rearranged and transformed to the following flexible form:
$\sum_{h<\exp(8/\varepsilon^{2})}\sum_{d|P(z)}\lambda^{+}(d,h)r\big(\mathcal{A},d\big),$
where the coefficients $\lambda^{+}(d,h)$ satisfy $\big|\lambda_{h}^{+}(d,h)\big|\leq 1$ and vanish for $d>D$. Especially, $\lambda^{+}(d,h)$ are well factorable of level $D$. Then by applying Lemma 2.2 with $D=x^{4/7-\varepsilon}$ and $L_{1}=1$, $L_{2}=x^{1-c}$, we obtain for the error term
$\sum_{\begin{subarray}{c}d|P(z)\end{subarray}}\lambda^{+}_{d}\,r(\mathcal{A},d)$ $=\sum_{h<\exp(8/\varepsilon^{2})}\sum_{\begin{subarray}{c}d\leq x^{4/7-\varepsilon}\end{subarray}}\lambda^{+}(d,h)r\big(\mathcal{A},d\big)$
$=\sum_{h<\exp(8/\varepsilon^{2})}\sum_{\begin{subarray}{c}d\leq x^{4/7-\varepsilon}\end{subarray}}\lambda^{+}(d,h)\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,(\ell,d)=1\end{subarray}}\bigg(\pi(x;\ell,-1,d)-\frac{\pi(x/\ell)}{\varphi(d)}\bigg)$
$\ll_{\varepsilon}\frac{x}{(\log x)^{A}}$ (3.3)
for any $A>0$, which is negligible.
Now it remains to estimate the main term $M(\mathcal{A},\mathcal{P},z)$. Changing the order of summation, we find that
$M(\mathcal{A},\mathcal{P},z)=$ $\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell\end{subarray}}\pi(x/\ell)\sum_{\begin{subarray}{c}d|P(z)\\ (d,\,\ell)=1\end{subarray}}\frac{\lambda^{+}_{d}}{\varphi(d)}$
$=$ $\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell\end{subarray}}\pi(x/\ell)\sum_{\begin{subarray}{c}d|P^{\prime}(z)\end{subarray}}\frac{\lambda^{+}_{d}}{\varphi(d)}$
@@
$P^{\prime}(z)=\prod_{p\leq z,\,p\in\mathscr{P}}p,\quad\mathscr{P}=\big\{p:\ p\nmid\ell\big\}.$
Then we apply Lemma 2.1 with $\omega(p)=p/\varphi(p),\,z=D=x^{4/7-\varepsilon}$ and some large constant $L$ getting
$M(\mathcal{A},\mathcal{P},z)$ $\leq\Big\{F(1)+o(1)\Big\}\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell\end{subarray}}\frac{x}{\ell\log(x/\ell)}\ \prod_{\begin{subarray}{c}2<p\leq z\\ p\nmid\ell\end{subarray}}\Big(1-\frac{1}{p-1}\Big)$
$=\Big\{F(1)+o(1)\Big\}\,\prod_{\begin{subarray}{c}2<p\leq z\end{subarray}}\Big(\frac{p-2}{p-1}\Big)\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell\end{subarray}}\frac{x}{\ell\log(x/\ell)}\prod_{2<p\leq z,\,p|\ell}\Big(\frac{p-1}{p-2}\Big)$
$=\Big\{F(1)+o(1)\Big\}\,\prod_{\begin{subarray}{c}2<p\leq z\end{subarray}}\Big(\frac{p-2}{p-1}\Big)\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\end{subarray}}\frac{x}{2\ell\log(x/2\ell)}\prod_{2<p\leq z,\,p|2\ell}\Big(\frac{p-1}{p-2}\Big)$
Considering
$\frac{1}{\log(x/2\ell)}=\frac{1}{\log(x/\ell)}\Big\{1+O\Big(\frac{1}{\log x}\Big)\Big\}\quad{\rm and}\quad\prod_{2<p\leq z,\,p|\ell}\Big(\frac{p-1}{p-2}\Big)=\prod_{p>2,\,p|\ell}\Big(\frac{p-1}{p-2}\Big)$
since $z=x^{4/7-\varepsilon}>x^{1-c}\geq\ell$, we then arrive at
$M(\mathcal{A},\mathcal{P},z)$ $=x\Big\{\frac{F(1)}{2}+o(1)\Big\}\,\prod_{\begin{subarray}{c}2<p\leq z\end{subarray}}\Big(\frac{p-2}{p-1}\Big)\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\end{subarray}}\frac{1}{\ell\log(x/\ell)}\prod_{p>2,\,p|\ell}\Big(\frac{p-1}{p-2}\Big).$ (3.4)
Next we define the multiplicative function $H(\ell)$ by
$H(\ell):=\prod_{p>2,\,p|\ell}\Big(\frac{p-1}{p-2}\Big).$
The partial sum of $H(\ell)$ is well studied by La Bretèche, Pomerance and Tenenbaum [9] by employing the Selberg-Delange method: for any nonnegative integer $v$,
$\sum_{\begin{subarray}{c}\ell\leq y\\ (\ell,\,a)=1\end{subarray}}H(2^{v}\ell)\sim\frac{Ay}{2^{\varepsilon(a)}H(a)}\qquad(y\to\infty),$
where $\varepsilon(a)=1$ if $2\mid a$ and $\varepsilon(a)=0$ if $2\nmid a$, and $A$ is an absolute constant defined by
$A:=\prod_{p>2}\Big(1+\frac{1}{p(p-2)}\Big).$
By taking $a=1$ in the above asymptotic formula and by partial summation, we infer that
$\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\end{subarray}}\frac{1}{\ell\log(x/\ell)}H(\ell)=\int^{x^{1-c}}_{1}\frac{{A\rm{d}}t}{t\log(x/t)}+o(1)=A\log\frac{1}{c}+o(1).$ (3.5)
Inserting (3.5) into (3.6) we get
$M(\mathcal{A},\mathcal{P},z)$ $\leq x\Big\{\frac{F(1)}{2}+o(1)\Big\}A\log\frac{1}{c}\,\prod_{\begin{subarray}{c}2<p\leq z\end{subarray}}\Big(\frac{p-2}{p-1}\Big).$
Note that $F(1)=2\textrm{e}^{\gamma}$, and by Merten’s theorem
$\prod_{\begin{subarray}{c}2<p\leq z\end{subarray}}\Big(\frac{p-2}{p-1}\Big)$ $=\prod_{\begin{subarray}{c}2<p\leq x^{4/7-\varepsilon}\end{subarray}}\Big(\frac{p-2}{p-1}\Big)\Big(1-\frac{1}{p}\Big)^{-1}\prod_{\begin{subarray}{c}2<p\leq x^{4/7-\varepsilon}\end{subarray}}\Big(1-\frac{1}{p}\Big)$
$=\frac{2\textrm{e}^{-\gamma}+o(1)}{\log x^{4/7-\varepsilon}}\prod_{\begin{subarray}{c}2<p\leq x^{4/7-\varepsilon}\end{subarray}}\frac{p(p-2)}{(p-1)^{2}}$
$=\frac{\frac{7}{2}\textrm{e}^{-\gamma}+o(1)}{\log x}\prod_{\begin{subarray}{c}p>2\end{subarray}}\Big(1+\frac{1}{p(p-2)}\Big)^{-1}\Big\{1+O\Big(\frac{1}{x}\Big)\Big\}$
$=\frac{1}{\log x}\Big(\frac{7}{2}\textrm{e}^{-\gamma}A^{-1}+o(1)\Big),$
therefore we find
$M(\mathcal{A},\mathcal{P},z)$ $\leq\Big(-\frac{7}{2}\log c+o(1)\Big)\frac{x}{\log x}.$ (3.6)
Finally, combining (3), (3) and (3.6), we conclude that
$S(\mathcal{A},\mathcal{P},z)\leq\Big(-\frac{7}{2}\log c+o(1)\Big)\frac{x}{\log x},$
and hence, $T^{\prime}_{c}(x)$ is estimated from (3.1) as follows